OpenStudy (anonymous):

solve the equation using the zero-product property -9n(5n-5)=0

OpenStudy (thomas5267):

Zero-product property means that if a×b=0, then a=0, b=0 or a=b=0.

OpenStudy (anonymous):

how do i find the answer? i dont know how to do this problem @thomas5267

OpenStudy (thomas5267):

\[ -9n(5n-5)=0\\ -9n=0 \text{ or }5n-5=0 \] Solve it.

OpenStudy (zaraxx):

0,1 in case anyone else needs it